OpenStudy (aaaa):

The area of a triangle is 15m9n9 square units. If the base of the triangle is 5m5n6 units, what is the height of the triangle? A. 3m4n2 units B. 6m4n3 units C. 3m4n3 units D. 6m3n3 units

OpenStudy (anonymous):

look, ill explain it from scratch

OpenStudy (cwrw238):

area = 1/2 * base * height

OpenStudy (cwrw238):

height =[ 2 * area] / base

OpenStudy (cwrw238):

= 2 *15m9n9 / 5m5n6 = 6 m4n3